# ClinKD: Cross-Modal Clinical Knowledge Distiller For Multi-Task Medical Images

### v3

@@ 3.3 Reflective Correction Training
After the model distillation, we restart two rounds of supervised fine-tuning using the whole dataset. During the supervised fine-tuning process, the model’s outputs are recorded and compared with the ground truth. If the semantic similarity between the output and the ground truth falls below 80%, GPT-4o will correct the output based on the ground truth (retaining only the original meaning and anchor box coordinates) to enrich the diversity of the training set labels. The corrected data is then fed back into the training process for further learning.
@@ 3.4 Semantic-Aware Selective Generation
The responses generated by large language models exhibit randomness, and relying on a single inference may result in missing the optimal answer. Semantic-Aware Selective Generation means that instead of directly using the model’s single-round generation for evaluation, we allow the model to generate multiple samples. We then utilize the CLIP models to score the alignment between the generated text and image. Finally, we reorder the outputs based on these scores and select the one with the highest score as the final result. The pseudo code is shown in Algorithm 1.
Algorithm 1 Semantic-Aware Selective Generation
1:  Input: Image $I$, Model $M$
2:  Output: Optimal generated text $T_{final}$
3:  $T_{1},T_{2},\dots,T_{n}\leftarrow\text{GenerateSamples}(M,I)$
4:  for $i=1$ to $n$ do
5:     $S_{i}\leftarrow\text{CLIPScore}(T_{i},I)$
6:  end for
7:  $T_{ordered}\leftarrow\text{Sort}(T_{1},T_{2},\dots,T_{n})$
8:  $T_{final}\leftarrow T_{ordered}[1]$
9:  return  $T_{final}$
@@ 4 Experiments
LLaVA-Med [18] Med-GRIT-Test30k [39, 7] 0 2.75 8.18 11.20 5.53
BiRD [14] Med-GRIT-Test30k [39, 7] 53.92 65.33 55.23 52.17 56.66
ClinKD (Ours) Med-GRIT-Test30k [39, 7] 67.51 82.35 70.56 65.69 71.53
LLaVA-Med [18] LLaVA-Med-qa0.2k [46] - - - 20.04 -
BiRD [14] LLaVA-Med-qa0.2k [46] - - - 10.55 -
ClinKD (Ours) LLaVA-Med-qa0.2k [46] - - - 13.24 -
Table 1: Comparison with LLaVA-Med [18] and BiRD (previous SOTA) [14]. The best scores are shown in bold.
@@
Qwen2-VL [34] 55.43 (+1.51) 68.33 (+3.00) 57.39 (+2.16) 62.23 (+10.06) 61.37 (+4.71)
BiRD [14] 53.92 65.33 55.23 52.17 56.66
@@
Table 2: Ablation study for different models and proposed methods. The value which is blue shows how much the metric scores improved. The best scores are shown in bold. The MCG-RoPE, Pseudo-KD and SASG stand for Med-CLIP Guided RoPE, Pseudo-Labels Medical Knowledge Distillation and Semantic-Aware Selective Generation respectively.
@@
BiRD [14] 44.47$\pm$0.14 29.26$\pm$0.11 41.73$\pm$0.16 56.46$\pm$0.15 53.60$\pm$0.12 75.63$\pm$0.17 84.15$\pm$0.13 46.04$\pm$0.18 53.92$\pm$0.14
BiRD [14] + RoPE-Mixed [12] 46.90$\pm$0.16 53.11$\pm$0.13 42.03$\pm$0.17 60.84$\pm$0.14 53.49$\pm$0.15 78.25$\pm$0.11 85.16$\pm$0.12 52.83$\pm$0.13 59.08$\pm$0.19
BiRD [14] + MCG-RoPE (Ours) 51.22$\pm$0.14 49.40$\pm$0.18 45.92$\pm$0.17 68.60$\pm$0.13 60.37$\pm$0.19 88.28$\pm$0.12 86.59$\pm$0.15 61.88$\pm$0.16 64.03$\pm$0.15
BiRD [14] + $V_{k}D$  [30] 58.77$\pm$0.12 50.95$\pm$0.16 45.69$\pm$0.11 65.89$\pm$0.18 57.63$\pm$0.14 93.32$\pm$0.15 79.27$\pm$0.12 59.90$\pm$0.15 63.93$\pm$0.19
BiRD [14] + Pseudo-KD (Ours) 62.59$\pm$0.15 52.86$\pm$0.17 47.19$\pm$0.13 66.28$\pm$0.16 59.84$\pm$0.11 93.68$\pm$0.18 85.37$\pm$0.14 62.87$\pm$0.19 66.33$\pm$0.13
@@
BiRD [14] 34.76$\pm$0.11 61.79$\pm$0.15 53.74$\pm$0.14 - 60.40$\pm$0.17 96.61$\pm$0.18 - 84.65$\pm$0.12 65.33$\pm$0.14
BiRD [14] + RoPE-Mixed [12] 39.78$\pm$0.13 62.13$\pm$0.19 74.27$\pm$0.11 - 61.33$\pm$0.13 96.61$\pm$0.17 - 84.65$\pm$0.15 69.80$\pm$0.12
BiRD [14] + MCG-RoPE (Ours) 44.47$\pm$0.15 64.20$\pm$0.18 83.75$\pm$0.11 - 62.19$\pm$0.14 96.61$\pm$0.19 - 85.22$\pm$0.16 72.74$\pm$0.13
BiRD [14] + $V_{k}D$ [30] 53.19$\pm$0.17 69.10$\pm$0.14 90.84$\pm$0.19 - 61.71$\pm$0.10 100$\pm$0.00 - 81.15$\pm$0.12 76.00$\pm$0.15
BiRD [14] + Pseudo-KD (Ours) 58.47$\pm$0.16 69.92$\pm$0.12 95.83$\pm$0.18 - 66.32$\pm$0.11 100$\pm$0.00 - 86.58$\pm$0.13 79.52$\pm$0.14
@@
BiRD [14] 41.88$\pm$0.12 51.69$\pm$0.14 37.39$\pm$0.18 47.95$\pm$0.15 54.07$\pm$0.10 77.44$\pm$0.16 48.73$\pm$0.13 82.65$\pm$0.19 55.23$\pm$0.14
BiRD [14] + RoPE-Mixed [12] 43.51$\pm$0.15 51.88$\pm$0.18 54.61$\pm$0.13 52.83$\pm$0.19 54.07$\pm$0.11 68.15$\pm$0.17 53.99$\pm$0.12 65.28$\pm$0.14 55.54$\pm$0.13
BiRD [14] + MCG-RoPE (Ours) 50.73$\pm$0.14 53.11$\pm$0.16 60.33$\pm$0.12 70.16$\pm$0.18 54.07$\pm$0.11 70.18$\pm$0.19 61.52$\pm$0.15 70.46$\pm$0.17 61.32$\pm$0.13
BiRD [14] + $V_{k}D$  [30] 51.33$\pm$0.13 54.39$\pm$0.17 62.83$\pm$0.19 74.31$\pm$0.14 54.07$\pm$0.15 76.23$\pm$0.12 67.89$\pm$0.18 72.15$\pm$0.11 64.15$\pm$0.16
BiRD [14] + Pseudo-KD (Ours) 52.78$\pm$0.12 58.41$\pm$0.15 70.39$\pm$0.18 75.85$\pm$0.13 55.22$\pm$0.16 81.34$\pm$0.11 71.54$\pm$0.17 72.15$\pm$0.14 67.21$\pm$0.19
RC SPICE [2] ClinKD (Ours) 53.94$\pm$0.14 58.41$\pm$0.12 71.47$\pm$0.18 79.91$\pm$0.15 52.10$\pm$0.11 83.92$\pm$0.19 84.51$\pm$0.13 80.19$\pm$0.16 70.56$\pm$0.17
BiRD [14] 47.01$\pm$0.14 49.35$\pm$0.18 37.17$\pm$0.15 57.15$\pm$0.12 39.91$\pm$0.17 72.13$\pm$0.13 48.87$\pm$0.16 65.78$\pm$0.19 52.17$\pm$0.11
BiRD [14] + RoPE-Mixed [12] 57.10$\pm$0.17 50.89$\pm$0.13 42.55$\pm$0.16 69.28$\pm$0.11 41.63$\pm$0.18 73.32$\pm$0.14 59.63$\pm$0.15 61.80$\pm$0.10 57.03$\pm$0.19
BiRD [14] + MCG-RoPE (Ours) 60.53$\pm$0.16 53.77$\pm$0.11 54.10$\pm$0.18 62.01$\pm$0.12 43.88$\pm$0.15 75.33$\pm$0.14 62.89$\pm$0.19 64.53$\pm$0.13 59.63$\pm$0.17
BiRD [14] + $V_{k}D$  [30] 62.33$\pm$0.10 53.77$\pm$0.15 52.27$\pm$0.19 62.01$\pm$0.16 47.29$\pm$0.13 75.33$\pm$0.17 68.04$\pm$0.12 70.31$\pm$0.18 61.42$\pm$0.14
BiRD [14] + Pseudo-KD (Ours) 66.85$\pm$0.11 55.73$\pm$0.14 52.27$\pm$0.17 64.74$\pm$0.15 50.58$\pm$0.12 78.21$\pm$0.19 70.02$\pm$0.13 76.16$\pm$0.18 64.32$\pm$0.16
MIA mBMR [14] ClinKD (Ours) 66.85$\pm$0.13 56.82$\pm$0.16 54.01$\pm$0.18 67.08$\pm$0.14 50.58$\pm$0.11 81.00$\pm$0.15 70.02 $\pm$0.17 76.05$\pm$0.19 65.69$\pm$0.12
BiRD [14] 43.03$\pm$0.12 48.02$\pm$0.11 42.51$\pm$0.17 53.85$\pm$0.19 51.99$\pm$0.14 80.45$\pm$0.16 60.58$\pm$0.13 69.78$\pm$0.15 -
BiRD [14] + RoPE-Mixed [12] 46.82$\pm$0.17 54.50$\pm$0.12 53.37$\pm$0.15 60.98$\pm$0.13 52.63$\pm$0.18 79.08$\pm$0.10 66.26$\pm$0.19 66.14$\pm$0.11 -
BiRD [14] + MCG-RoPE (Ours) 51.74$\pm$0.13 55.12$\pm$0.19 61.03$\pm$0.12 66.92$\pm$0.17 55.13$\pm$0.14 82.60$\pm$0.15 70.33$\pm$0.11 70.52$\pm$0.18 -
BiRD [14] + $V_{k}D$ [30] 56.41$\pm$0.16 57.05$\pm$0.11 62.91$\pm$0.19 67.40$\pm$0.14 55.18$\pm$0.18 86.22$\pm$0.10 71.73$\pm$0.15 70.88$\pm$0.13 -
BiRD [14] + Pseudo-KD (Ours) 60.17$\pm$0.14 59.23$\pm$0.16 66.42$\pm$0.12 68.96$\pm$0.19 57.99$\pm$0.13 88.31$\pm$0.15 75.64$\pm$0.17 74.44$\pm$0.10 -
@@
Table 3: Comparison with other SOTA methods on Med-GRIT-Test30k [39, 7]. We evaluate the performance on eight modalities.
4.1 Datasets
Med-GRIT-270k [14]. The Med-GRIT-270k dataset comprises 270k samples, with 240k allocated for training and the remaining 30k for multi-task evaluation. These samples are collected from eight medical modalities: CT, MR, X-ray, PET, Endoscopy, Dermoscopy, Fundus and Ultrasound. Additionally, we randomly select 30K samples from the entire dataset for pseudo-label generation, which is subsequently used by the Pseudo-Labels Medical Knowledge Distillation. This fine-grained multi-task dataset not only demands that the model accurately understands medical images, but also requires it to precisely classify and ground pathological regions and physiological structures. This poses a significant challenge for MLLMs.
LLaVA-Med-qa0.2k [46]. The LLaVA-Med-qa0.2k [46] contains 0.2k QA pairs of MIA task which are used for evaluating the models’ ability to analyze medical images. It consists of 50 images with caption pairs and two types of questions: conversation and detailed description.
The images are not only from common modalities such as CT and MR, but also from some modalities that are not available in the Med-GRIT-270k [14] dataset, such as 3D reconstruction. The wide range of data sources evaluates the capabilities of MLLMs in general medical scenarios.
4.2 Evaluation Metrics
We use Recall@0.5 for Visual Grounding (VG), Recall for Referring Object Classification (ROC), SPICE [2] for Referring Caption (RC), and mBMR [14] for Medical Image Analysis (MIA). The final overall result is computed as the mean of the evaluation scores across these four tasks.
For the evaluation on LLaVA-Med-qa0.2k [18] which has 0.2k QA pairs of MIA task, we use mBMR as the metric.
4.3 Implementation Details
We use Qwen2-VL [34] as the base model for ClinKD. For the Med-CLIP Guided RoPE component of ClinKD, in order to ensure that the position index is an integer, we resample the image resolution to $w=h$. During the distillation process, we set $\alpha=0.5$, $T=0.5$. The number of distillation epochs is 2, with an additional 2 epochs for subsequent training. The entire training process is conducted using 6 NVIDIA A100 40G GPUs. The learning rate is set to $2\times 10^{-5}$, and the AdamW optimizer is employed. The learning scheduler follows a cosine schedule. For model generation, we use non-sampling beam search with a beam size of 3.
4.4 Analysis
4.4.1 Comparison with the State-of-the-Art Methods
Evaluation on the Med-GRIT-Test30k [39, 7]. As shown in Table 1, we compare ClinKD with BiRD [14] and LLaVA-Med [18]. Results show that ClinKD achieves scores of 67.51%, 82.35%, 70.56%, and 65.69% in the VG, ROC, RC, and MIA tasks respectively, outperforming other comparisons by a clear margin in all tasks. On average, ClinKD outperforms BiRD [14] and LLaVA-Med [18] by 14.87% and 66.00% respectively. ClinKD’s advantages lie in its more adequate image-text alignment and the prior medical knowledge provided by Pseudo-Labels Medical Knowledge Distillation.
Evaluation on the LLaVA-Med-qa0.2k [46]. Table 1 represents the models’ performance on LLaVA-Med-qa0.2k [46]. LLaVA-Med [18] still achieves the best performance. Although ClinKD has not yet surpassed LLaVA-Med [18] on this dataset, it still achieved a 2.69% improvement over BiRD [14] in the MIA task. The limitation may be caused by the different sources of medical data so that the ClinKD may fail to leverage medical knowledge to analyze medical images.
Figure 4: Effect of value of $\alpha$ in Pseudo-Labels Medical Knowledge Distillation.
4.4.2 Ablation study
Effect of Qwen2-VL [34]. Table 2 presents the ablation study of proposed methods. To investigate the impact of model upgrade on the experiment, we applied the same training strategy to Qwen2-VL as we did to BiRD [14]. The results show that the model upgrade only has a significant effect on the MIA task, which shows Qwen2-VL [34] outperforms BiRD by 10.06%.
Effect of Pseudo-Labels Medical Knowledge Distillation. Table 2 represents improvements achieved by using Pseudo-Labels Medical Knowledge Distillation that leads to enhancements by 12.41%, 14.19%, 11.98% and 12.15% in the VG, ROC, RC and MIA tasks, respectively. The results illustrate that providing prior medical knowledge by using Pseudo-Labels Medical Knowledge Distillation can fill MLLMs’ gap of medical knowledge, boosting capabilities of adapting to new medical knowledge.
Figure 4 shows the effect of different values of $\alpha$ in Eq. (16). In this experiment, the average performance achieves the best when we set $\alpha=0.5$. Both excessively high and low alpha values can lead to performance degradation. This might be because a low $\alpha$ value prevents ClinKD from learning sufficient prior medical knowledge, while an overly high alpha value tends to make ClinKD generate irrational pseudo-labels that contradict common medical knowledge.

### v4

@@ 3.3 Reflective Correction Training
In this part, during each round of supervised fine-tuning, we record the model’s prediction $\hat{\mathbf{y}}=(\mathbf{t}_{o},\mathbf{c})$ and the corresponding ground-truth label $\mathbf{y}=(\mathbf{t}_{gt},\mathbf{c})$. We then compute a weighted cosine similarity to assess semantic alignment and, if necessary, invoke GPT-4o to “correct” the output in order to enrich label diversity. The procedure is as follows:
First, map the text labels into a $d$-dimensional embedding space to obtain $\mathbf{e}_{o},\mathbf{e}_{gt}\in\mathbb{R}^{d}$, while retaining the shared anchor-box coordinate vector $\mathbf{c}\in\mathbb{R}^{4}$.
Next, define the weighted cosine similarity as
$S(\mathbf{e}_{o},\mathbf{e}_{gt};\mathbf{c})=\frac{\mu\,\langle\mathbf{e}_{o},\mathbf{e}_{gt}\rangle+\nu\,\langle\mathbf{c},\mathbf{c}\rangle}{\sqrt{\bigl{(}\mu\,\|\mathbf{e}_{o}\|^{2}+\nu\,\|\mathbf{c}\|^{2}\bigr{)}\,\bigl{(}\mu\,\|\mathbf{e}_{gt}\|^{2}+\nu\,\|\mathbf{c}\|^{2}\bigr{)}}}\,,$ (15)
[…]
We then compare $S$ against $\tau$ and decide whether to trigger GPT-4o correction:
$\hat{\mathbf{y}}^{\prime}=\begin{cases}\mathrm{GPT4o\_Correct}\bigl{(}\hat{\mathbf{y}},\,\mathbf{y}\bigr{)},&S(\mathbf{e}_{o},\mathbf{e}_{gt};\mathbf{c})<\tau,\\[6.0pt] \hat{\mathbf{y}},&S(\mathbf{e}_{o},\mathbf{e}_{gt};\mathbf{c})\geq\tau,\end{cases}$ (16)
[…]
@@ 3.4 Semantic-Aware Selective Generation
The responses generated by large language models exhibit randomness, and relying on a single inference may result in missing the optimal answer. Semantic-Aware Selective Generation means that instead of directly using the model’s single-round generation for evaluation, we allow the model to generate multiple samples. We then utilize the CLIP models to score the alignment between the generated text and image. Finally, we reorder the outputs based on these scores and select the one with the highest score as the final result. The pseudo code is shown in Algorithm.
@@ 4 Experiments
4.1 Datasets
We train ClinKD on the Med-GRIT-270k huang2024BiRD dataset, and evaluate the performance on Med-GRIT-30k and LLaVA-Med-qa0.2k zhang2025biomedclipmultimodalbiomedicalfoundation dataset.
We alse test the few-shot performance on VQA-RAD lau2018dataset , SLAKE liu2021slake and Path-VQA he2020pathvqa . For more information, please refer to Appendix.
4.2 Evaluation Metrics
We use Recall@0.5 for Visual Grounding (VG), Recall for Referring Object Classification (ROC), SPICE anderson2016spicesemanticpropositionalimage for Referring Caption (RC), and mBMR huang2024BiRD for Medical Image Analysis (MIA). The final overall result is computed as the mean of the evaluation scores across these four tasks.
For the evaluation on LLaVA-Med-qa0.2k li2024llava which has 0.2k QA pairs of MIA task, we use mBMR as the metric.
In few-shot experiments, we use BLEU-1 and F1-score to evaluate models’ performance.
4.3 Analysis
4.3.1 Comparison with the State-of-the-Art Methods
Evaluation on the Med-GRIT-Test30k ye2023samed2d20m ; cheng2023sammed2d . As shown in Table 1, we compare ClinKD with BiRD huang2024BiRD and LLaVA-Med li2024llava . Results show that ClinKD achieves scores of 67.51%, 82.35%, 70.56%, and 65.69% in the VG, ROC, RC, and MIA tasks respectively, outperforming other comparisons by a clear margin in all tasks. On average, ClinKD outperforms BiRD huang2024BiRD and LLaVA-Med li2024llava by 14.87% and 66.00% respectively. ClinKD’s advantages lie in its more adequate image-text alignment and the prior medical knowledge provided by Pseudo-KD.
Evaluation on the LLaVA-Med-qa0.2k zhang2025biomedclipmultimodalbiomedicalfoundation . Table 1 represents the models’ performance on LLaVA-Med-qa0.2k zhang2025biomedclipmultimodalbiomedicalfoundation . Our ClinKD model demonstrates superior performance on this dataset, outperforming both LLaVA-Med (by 3.50%) and BiRD (by 2.48%). The relatively low mBMR scores can be attributed to the notable language style discrepancies observed between the training and testing datasets.
@@ 4.3.1 Comparison with the State-of-the-Art Methods
LLaVA-Med li2024llava Med-GRIT-Test30k ye2023samed2d20m ; cheng2023sammed2d 0 2.75 8.18 11.20 5.53
BiRD huang2024BiRD Med-GRIT-Test30k ye2023samed2d20m ; cheng2023sammed2d 53.92 65.33 55.23 52.17 56.66
ClinKD (Ours) Med-GRIT-Test30k ye2023samed2d20m ; cheng2023sammed2d 67.51 82.35 70.56 65.69 71.53
LLaVA-Med li2024llava LLaVA-Med-qa0.2k zhang2025biomedclipmultimodalbiomedicalfoundation - - - 20.04 -
BiRD huang2024BiRD LLaVA-Med-qa0.2k zhang2025biomedclipmultimodalbiomedicalfoundation - - - 21.06 -
ClinKD (Ours) LLaVA-Med-qa0.2k zhang2025biomedclipmultimodalbiomedicalfoundation - - - 23.54 -
Table 1: Comparison with LLaVA-Med li2024llava and BiRD (previous SOTA) huang2024BiRD . The best scores are shown in bold.
@@
Qwen2-VL wang2024qwen2vlenhancingvisionlanguagemodels 55.43 (+1.51) 68.33 (+3.00) 57.39 (+2.16) 62.23 (+10.06) 61.37 (+4.71)
BiRD huang2024BiRD 53.92 65.33 55.23 52.17 56.66
@@
Table 2: Ablation study for different models and proposed methods. The value which is blue shows how much the metric scores improved. The best scores are shown in bold. The MCG-RoPE, Pseudo-KD and SASG stand for Med-CLIP Guided RoPE, Adaptive Confidence-Margin Curriculum Pseudo-KD and Semantic-Aware Selective Generation respectively.
4.3.2 Ablation study
Effect of Qwen2-VL wang2024qwen2vlenhancingvisionlanguagemodels . Table 2 presents the ablation study of proposed methods. To investigate the impact of model upgrade on the experiment, we applied the same training strategy to Qwen2-VL as we did to BiRD huang2024BiRD . The results show that the model upgrade only has a significant effect on the MIA task, which shows Qwen2-VL wang2024qwen2vlenhancingvisionlanguagemodels outperforms BiRD by 10.06%.
Effect of Adaptive Confidence-Margin Curriculum Pseudo-KD. Table 2 represents improvements achieved by using Pseudo-KD that leads to enhancements by 12.41%, 14.19%, 11.98% and 12.15% in the VG, ROC, RC and MIA tasks, respectively. The results illustrate that providing prior medical knowledge by using Pseudo-KD can fill MLLMs’ gap of medical knowledge, boosting capabilities of adapting to new medical knowledge. Figure shows the effect of different values of $\alpha$ in Eq. (14). In this experiment, the average performance achieves the best when we set $\alpha=0.5$. Both excessively high and low alpha values can lead to performance degradation. This might be because a low $\alpha$ value prevents ClinKD from learning sufficient prior medical knowledge, while an overly high alpha value tends to make ClinKD generate irrational pseudo-labels that contradict common medical knowledge.
@@ 4.3.2 Ablation study
BiRD huang2024BiRD 44.47$\pm$0.14 29.26$\pm$0.11 41.73$\pm$0.16 56.46$\pm$0.15 53.60$\pm$0.12 75.63$\pm$0.17 84.15$\pm$0.13 46.04$\pm$0.18 53.92$\pm$0.14
BiRD huang2024BiRD + RoPE-Mixed heo2024rotarypositionembeddingvision 46.90$\pm$0.16 53.11$\pm$0.13 42.03$\pm$0.17 60.84$\pm$0.14 53.49$\pm$0.15 78.25$\pm$0.11 85.16$\pm$0.12 52.83$\pm$0.13 59.08$\pm$0.19
BiRD huang2024BiRD + MCG-RoPE (Ours) 51.22$\pm$0.14 49.40$\pm$0.18 45.92$\pm$0.17 68.60$\pm$0.13 60.37$\pm$0.19 88.28$\pm$0.12 86.59$\pm$0.15 61.88$\pm$0.16 64.03$\pm$0.15
BiRD huang2024BiRD + $V_{k}D$  Miles_2024_CVPR 58.77$\pm$0.12 50.95$\pm$0.16 45.69$\pm$0.11 65.89$\pm$0.18 57.63$\pm$0.14 93.32$\pm$0.15 79.27$\pm$0.12 59.90$\pm$0.15 63.93$\pm$0.19
BiRD huang2024BiRD + Pseudo-KD (Ours) 62.59$\pm$0.15 52.86$\pm$0.17 47.19$\pm$0.13 66.28$\pm$0.16 59.84$\pm$0.11 93.68$\pm$0.18 85.37$\pm$0.14 62.87$\pm$0.19 66.33$\pm$0.13
@@
BiRD huang2024BiRD 34.76$\pm$0.11 61.79$\pm$0.15 53.74$\pm$0.14 - 60.40$\pm$0.17 96.61$\pm$0.18 - 84.65$\pm$0.12 65.33$\pm$0.14
BiRD huang2024BiRD + RoPE-Mixed heo2024rotarypositionembeddingvision 39.78$\pm$0.13 62.13$\pm$0.19 74.27$\pm$0.11 - 61.33$\pm$0.13 96.61$\pm$0.17 - 84.65$\pm$0.15 69.80$\pm$0.12
BiRD huang2024BiRD + MCG-RoPE (Ours) 44.47$\pm$0.15 64.20$\pm$0.18 83.75$\pm$0.11 - 62.19$\pm$0.14 96.61$\pm$0.19 - 85.22$\pm$0.16 72.74$\pm$0.13
BiRD huang2024BiRD + $V_{k}D$ Miles_2024_CVPR 53.19$\pm$0.17 69.10$\pm$0.14 90.84$\pm$0.19 - 61.71$\pm$0.10 100$\pm$0.00 - 81.15$\pm$0.12 76.00$\pm$0.15
BiRD huang2024BiRD + Pseudo-KD (Ours) 58.47$\pm$0.16 69.92$\pm$0.12 95.83$\pm$0.18 - 66.32$\pm$0.11 100$\pm$0.00 - 86.58$\pm$0.13 79.52$\pm$0.14
@@
BiRD huang2024BiRD 41.88$\pm$0.12 51.69$\pm$0.14 37.39$\pm$0.18 47.95$\pm$0.15 54.07$\pm$0.10 77.44$\pm$0.16 48.73$\pm$0.13 82.65$\pm$0.19 55.23$\pm$0.14
BiRD huang2024BiRD + RoPE-Mixed heo2024rotarypositionembeddingvision 43.51$\pm$0.15 51.88$\pm$0.18 54.61$\pm$0.13 52.83$\pm$0.19 54.07$\pm$0.11 68.15$\pm$0.17 53.99$\pm$0.12 65.28$\pm$0.14 55.54$\pm$0.13
BiRD huang2024BiRD + MCG-RoPE (Ours) 50.73$\pm$0.14 53.11$\pm$0.16 60.33$\pm$0.12 70.16$\pm$0.18 54.07$\pm$0.11 70.18$\pm$0.19 61.52$\pm$0.15 70.46$\pm$0.17 61.32$\pm$0.13
BiRD huang2024BiRD + $V_{k}D$  Miles_2024_CVPR 51.33$\pm$0.13 54.39$\pm$0.17 62.83$\pm$0.19 74.31$\pm$0.14 54.07$\pm$0.15 76.23$\pm$0.12 67.89$\pm$0.18 72.15$\pm$0.11 64.15$\pm$0.16
BiRD huang2024BiRD + Pseudo-KD (Ours) 52.78$\pm$0.12 58.41$\pm$0.15 70.39$\pm$0.18 75.85$\pm$0.13 55.22$\pm$0.16 81.34$\pm$0.11 71.54$\pm$0.17 72.15$\pm$0.14 67.21$\pm$0.19
RC SPICE anderson2016spicesemanticpropositionalimage ClinKD (Ours) 53.94$\pm$0.14 58.41$\pm$0.12 71.47$\pm$0.18 79.91$\pm$0.15 52.10$\pm$0.11 83.92$\pm$0.19 84.51$\pm$0.13 80.19$\pm$0.16 70.56$\pm$0.17
BiRD huang2024BiRD 47.01$\pm$0.14 49.35$\pm$0.18 37.17$\pm$0.15 57.15$\pm$0.12 39.91$\pm$0.17 72.13$\pm$0.13 48.87$\pm$0.16 65.78$\pm$0.19 52.17$\pm$0.11
BiRD huang2024BiRD + RoPE-Mixed heo2024rotarypositionembeddingvision 57.10$\pm$0.17 50.89$\pm$0.13 42.55$\pm$0.16 69.28$\pm$0.11 41.63$\pm$0.18 73.32$\pm$0.14 59.63$\pm$0.15 61.80$\pm$0.10 57.03$\pm$0.19
BiRD huang2024BiRD + MCG-RoPE (Ours) 60.53$\pm$0.16 53.77$\pm$0.11 54.10$\pm$0.18 62.01$\pm$0.12 43.88$\pm$0.15 75.33$\pm$0.14 62.89$\pm$0.19 64.53$\pm$0.13 59.63$\pm$0.17
BiRD huang2024BiRD + $V_{k}D$  Miles_2024_CVPR 62.33$\pm$0.10 53.77$\pm$0.15 52.27$\pm$0.19 62.01$\pm$0.16 47.29$\pm$0.13 75.33$\pm$0.17 68.04$\pm$0.12 70.31$\pm$0.18 61.42$\pm$0.14
BiRD huang2024BiRD + Pseudo-KD (Ours) 66.85$\pm$0.11 55.73$\pm$0.14 52.27$\pm$0.17 64.74$\pm$0.15 50.58$\pm$0.12 78.21$\pm$0.19 70.02$\pm$0.13 76.16$\pm$0.18 64.32$\pm$0.16
MIA mBMR huang2024BiRD ClinKD (Ours) 66.85$\pm$0.13 56.82$\pm$0.16 54.01$\pm$0.18 67.08$\pm$0.14 50.58$\pm$0.11 81.00$\pm$0.15 70.02 $\pm$0.17 76.05$\pm$0.19 65.69$\pm$0.12
BiRD huang2024BiRD 43.03$\pm$0.12 48.02$\pm$0.11 42.51$\pm$0.17 53.85$\pm$0.19 51.99$\pm$0.14 80.45$\pm$0.16 60.58$\pm$0.13 69.78$\pm$0.15 -
BiRD huang2024BiRD + RoPE-Mixed heo2024rotarypositionembeddingvision 46.82$\pm$0.17 54.50$\pm$0.12 53.37$\pm$0.15 60.98$\pm$0.13 52.63$\pm$0.18 79.08$\pm$0.10 66.26$\pm$0.19 66.14$\pm$0.11 -
BiRD huang2024BiRD + MCG-RoPE (Ours) 51.74$\pm$0.13 55.12$\pm$0.19 61.03$\pm$0.12 66.92$\pm$0.17 55.13$\pm$0.14 82.60$\pm$0.15 70.33$\pm$0.11 70.52$\pm$0.18 -
BiRD huang2024BiRD + $V_{k}D$ Miles_2024_CVPR 56.41$\pm$0.16 57.05$\pm$0.11 62.91$\pm$0.19 67.40$\pm$0.14 55.18$\pm$0.18 86.22$\pm$0.10 71.73$\pm$0.15 70.88$\pm$0.13 -
BiRD huang2024BiRD + Pseudo-KD (Ours) 60.17$\pm$0.14 59.23$\pm$0.16 66.42$\pm$0.12 68.96$\pm$0.19 57.99$\pm$0.13 88.31$\pm$0.15 75.64$\pm$0.17 74.44$\pm$0.10 -
@@
Table 3: Comparison with other SOTA methods on Med-GRIT-Test30k ye2023samed2d20m ; cheng2023sammed2d . We evaluate the performance on eight modalities.